x+y=10
y+z=5
z+x=4
solve by Cramers rule

Answer by richwmiller(17219)

1x+1y+0*z=10
0*x+1y+1z=5
1x+0*y+1z=4
1,1,0,10
0,1,1,5
1,0,1,4
1,1,0
0,1,1
1,0,1
working on determinant 1
add down (0/1) *row 1 to row 2
0,1,1
1,1,0
0,1,1
1,0,1
working on determinant 1
add down (-1/1) *row 1 to row 3
0,-1,1
1,1,0
0,1,1
0,-1,1
working on determinant 1
divide row 2 by 1/1
0,1,1
1,1,0
0,1,1
0,-1,1
working on determinant 1
add down (1/1) *row 2 to row 3
0,0,2
1,1,0
0,1,1
0,0,2
working on determinant 1
divide row 3 by 2/1
0,0,1
1,1,0
0,1,1
0,0,1
1 1/1
1 1/1
2 2/1
det 2
# 1
determinant
1,1,10
0,1,5
1,0,4
working on determinant 2
divide row 1 by 1/1
1,1,10
1,1,10
0,1,5
1,0,4
working on determinant 2
add down (0/1) *row 1 to row 2
0,1,5
1,1,10
0,1,5
1,0,4
working on determinant 2
add down (-1/1) *row 1 to row 3
0,-1,-6
1,1,10
0,1,5
0,-1,-6
working on determinant 2
divide row 2 by 1/1
0,1,5
1,1,10
0,1,5
0,-1,-6
working on determinant 2
add down (1/1) *row 2 to row 3
0,0,-1
1,1,10
0,1,5
0,0,-1
working on determinant 2
divide row 3 by -1/1
0,0,1
1,1,10
0,1,5
0,0,1
1 1/1
1 1/1
-1 -1/1
det -1
# 2
-1/2=-0.5
z = -1/2
determinant
1,10,0
0,5,1
1,4,1
working on determinant 3
divide row 1 by 1/1
1,10,0
1,10,0
0,5,1
1,4,1
working on determinant 3
add down (0/1) *row 1 to row 2
0,5,1
1,10,0
0,5,1
1,4,1
working on determinant 3
add down (-1/1) *row 1 to row 3
0,-6,1
1,10,0
0,5,1
0,-6,1
working on determinant 3
divide row 2 by 5/1
0,1,1/5
1,10,0
0,1,1/5
0,-6,1
working on determinant 3
add down (6/1) *row 2 to row 3
0,0,11/5
1,10,0
0,1,1/5
0,0,11/5
working on determinant 3
divide row 3 by 11/5
0,0,1
1,10,0
0,1,1/5
0,0,1
1 1/1
5 5/1
2.2 11/5
det 11
# 3
11/2=5.5
y = 11/2
determinant
10,1,0
5,1,1
4,0,1
working on determinant 4
divide row 1 by 10/1
1,1/10,0
1,1/10,0
5,1,1
4,0,1
working on determinant 4
add down (-5/1) *row 1 to row 2
0,5/10,1
1,1/10,0
0,5/10,1
4,0,1
working on determinant 4
add down (-4/1) *row 1 to row 3
0,-4/10,1
1,1/10,0
0,5/10,1
0,-2/5,1
working on determinant 4
divide row 2 by 5/10
0,1,2
1,1/10,0
0,1,2
0,-2/5,1
working on determinant 4
add down (2/5) *row 2 to row 3
0,0,9/5
1,1/10,0
0,1,2
0,0,9/5
working on determinant 4
divide row 3 by 9/5
0,0,1
1,1/10,0
0,1,2
0,0,1
10 10/1
0.5 5/10
1.8 9/5
det 9
# 4
9/2=4.5
x = 9/2
Summary
determinant = 2
x = 9/2 = 4.5
y = 11/2 = 5.5
z = -1/2 = -0.5